# An average Brun–Titchmarsh theorem and shifted primes with a large prime factor

### v2

@@ 6. The final decomposition
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{301}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{304}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{305}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{306}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{307}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{308}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\tau}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{309}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{312}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
@@
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{401}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{402}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{403}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{404}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{405}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\tau}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{406}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{409}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
@@
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{501}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{502}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{503}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{504}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{505}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\tau}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{506}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{509}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$

### v3

@@ 6. The final decomposition
$\left(\int_{(t_{1},t_{2},t_{3})\in L_{301}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{304}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{305}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{306}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{307}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{308}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1})\in L_{309}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{312}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
@@
$\left(\int_{(t_{1},t_{2},t_{3})\in L_{401}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{402}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{403}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{404}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{405}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1})\in L_{406}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{409}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
@@
$\left(\int_{(t_{1},t_{2},t_{3})\in L_{501}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{502}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{503}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{504}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{505}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1})\in L_{506}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{509}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$